One CD player is 3 times as much as another. Together the two CD players would be $320. What is the price of the more expensive CD player?
Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

One CD player is 3 times as much as another. Together the two CD players would be $320. What is the price of the more expensive CD player?
-------------------
C = price of cheaper one
C + 3C = 320
-----------------
Did you make any effort?

Answer by jim_thompson5910(35256)

Let

x = price of cheaper player
y = price of more expensive player


Since "One CD player is 3 times as much as another", we can say (according to the definition made above)

y = 3x

Together the two CD players would be $320, so x and y must add to 320

x+y = 320

x + 3x = 320 .... plug in y = 3x

4x = 320

Solve for x. Then use it to find y. I'll let you finish up.